Solve the following equation for the variable. You must show your algebraic steps used to earn credit.

3x + 2 = 4
_______
3

Answer by stanbon(75887)

3x + 2 = 4
_______
3
-------------
cross-multiply to get:
3x+2 = 12
3x = 10
x = 10/3
=================
Cheers,
Stan H.